What does it mean to have a normal distribution of data?
The Normal or Gaussian distribution is a probability distribution which depends on two parameters: the mean and the variance (or standard deviation). In may real life situations measurements are found to be approximately normal. Furthermore, even if the underlying distribution of a variable is not normal, the mean of a number of repeated observations of the variable will approximate the normal distribution.
The term "approximate" is important because, although the heights of adult males (for example) appear to be normally distributed, the true normal distribution must allow negative heights whereas that is not physically possible!
It is 0.6050
1925 ways, 11C2*7C4
neutron moderation
What does it mean for an event to have a probability of 1?
Probability is on a scale of 1 to 0 and 1 is that .an event will certainly happen whereas 0 is that an event will not happen
It is 1 - prob(it does not rain today and tomorrow)= 1 - prob(it does not rain today)*prob(it does not rain tomorrow)
= 1 - [1- prob(it does rain today)]*[1 - prob(it does rain tomorrow)]
= 1 - [1 - 0.8]*[1 - 0.5]
= 1 - 0.2*0.5
= 1 - 0.1 = 0.9
The probability is 0. One card cannot be a club and a spade!
The Answers community requires more information for this question. Please edit your question to include more context. For example, telling us a fair what, without SHOUTING might me a good start.
What is the probability that the spinner will land 5?
The answer depends on the shape of the spinner and the numbers on it.
The probability is 1 and you do not need Matlab to get that answer - only a little bit of thought.
What is one of the four possible outcomes associated with a requested relief of liability?
Successful Recovery.
Number of samples = 42C4 = 42*41*40*39/24 = 111930
Which of the following is the best definition of sample space of a probability event?
Set of all possible outcomes.
How many 9 of spades in in a 52 pack of playing cards?
There is only one nine of spades in a standard deck of cards, so in probability, that would be 1/52
9/10
What is the probability of not being dealt a queen in a 52 card deck?
The answer will depend on the exact situation.If you are dealt a single card, the probability of that single card not being a queen is 12/13 - assuming you have no knowledge about the other cards.Here is another example. If you already hold three queens in your hand (and no other cards have been dealt), the probability of the next card being dealt being a queen is 1/49, so the probability of NOT getting a queen is 48/49 - higher than in the previous example.
Will Random sample always accurately predict the outcomes of an entire population?
No. Only a census can ACCURATELY predict the outcomes: a random sample cannot.
The probability is X/(X + 2W) where X is the number of orange gumballs.
How many black face cards are in a deck of 52 cards?
In a standard deck of 52 cards black face cards are the spades and claver.
There are 13 spade and 13 claver cards ( A,2,3,4,5,6,7,8,9,K,Q,J)
So there are a total of 26 black face cards in a deck of 52 cards.